# Exposing a Fatal Flaw in Sample-based Quantum Diagonalization Methods

### v1

@@ Abstract
Quantum Selected Configuration Interaction (QSCI) methods (also known as Sample-based Quantum Diagonalization, SQD) have emerged as promising near-term approaches to solving the electronic Schrödinger equation with quantum computers. In this work, we show that QSCI methods face fundamental limitations that severely hinder their practical applicability. Using the nitrogen molecule and the iron-sulfur cluster [2Fe-2S] as examples, we demonstrate that while QSCI can, in principle, yield high-quality CI expansions similar to classical SCI heuristics in some cases, the method struggles with inefficiencies in finding new determinants as sampling repeatedly selects already seen configurations. This inefficiency becomes especially pronounced when targeting high-accuracy results or sampling from an approximate ansatz. In cases where the sampling problem is not present, the resulting CI expansions are less compact than those generated from classical heuristics, rendering QSCI an overall more expensive method. Our findings suggest a fatal flaw in QSCI methods as the inescapable trade-off between finding sufficiently many determinants and generating compact, accurate CI expansions. This ultimately hinders utility in quantum chemistry applications as QSCI falls behind more efficient classical counterparts.
[…]
1 Introduction
@@
The full configuration interaction (FCI) method provides the exact solution to the electronic Schrödinger equation (within a given basis) and is thus considered the ground truth within quantum chemistry1, 2, 3. However, due to the exponential scaling in computing and memory requirements, exact FCI is limited to very small molecules. One of the largest demonstrations to date was by Gao et al. on propane (C3H8)/STO-3G4, who used a distributed FCI implementation on 256 nodes to treat a CI expansion of 1.3 trillion determinants. As a technical achievement, such work is certainly impressive. If one is willing to settle for near-exact FCI results (within, e.g., micro-Hartree precision), approximate methods such as selected configuration interaction (SCI) can obtain results much more affordably5. Various developments within electronic structure theory have focused on achieving near-FCI accuracy at a lower cost3, including methods based on the density matrix renormalization group (DMRG)6, many-body expanded FCI (MBE-FCI)7, 8, 9, and various flavors of SCI10, 11, 12, 13. Recently, such methods were put to the test in a “blind challenge” of predicting the ground-state energy of the benzene molecule (C6H6) in a cc-pVDZ basis set2. It revealed that a firm consensus between various near-exact FCI methods has still not been reached, even when agreement from calculations on smaller systems would have suggested so.
@@
Applying quantum computers within the field of quantum chemistry has received significant attention in recent years14, 15, 16, 17. This is partly due to the expectation that quantum computers should be a natural candidate for solving the electronic Schrödinger equation as they inherently operate within an exponentially large Hilbert space. In the long term, provided fault-tolerant devices become a reality, quantum computers could, for example, be used to obtain ground state energies using quantum phase estimation18, 19. However, current near-term quantum computers are noisy and rather limited in terms of qubit count, error rates, and execution speed, which limits their practical advantage over classical computing architectures20. This has motivated the development of hybrid algorithms that take advantage of both quantum and classical computing architectures, resulting in shallower quantum circuits, which are more realistic options for practical quantum computing applications in the near term21, 22, 23. The variational quantum eigensolver (VQE) algorithm24, 25, 26 has emerged as a particularly popular near-term method. VQE-type methods rely on a parameterized ansatz, which is iteratively optimized using a hybrid quantum-classical feedback loop. The quantum device prepares a state and measures the expectation values over Pauli operators. The classical computer collects these measurements to evaluate expectation values over the fermionic Hamiltonian and updates the parameters of the parameterized ansatz using classical optimization methods. Although popular, VQE-type methods have several serious drawbacks related to the measurement process27 (which inherently has statistical noise) and the expressibility of the ansatz28, 29, 26.
@@
The quantum selected configuration interaction (QSCI) method, introduced by Kanno et al.30, 31, addresses some of the challenges faced with VQE by drastically reducing the quantum workload, leveraging quantum devices solely for determinant selection. The central idea of the method is that some approximate ground state is prepared on a quantum computer, which is then used to draw samples corresponding to determinants in the computational basis. The CI Hamiltonian is then classically diagonalized in the basis of the selected determinants using conventional SCI diagonalization routines. We note that the classic diagonalization eventually becomes the computational bottleneck of the algorithm, provided enough determinants are discovered. Thus, the “quantum” part of QSCI lies solely in the selection of determinants for the SCI expansion. Such a division of work between quantum and classical resources has some clear advantages. First, since the quantum device is only used for selection, while energies are obtained classically, one removes the statistical errors associated with Hamiltonian measurement. Second, the CI vector is accessible on the classical device after the diagonalization, which allows computing expectation values over other quantum mechanical operators or correcting the energy of the variational wave function with Epstein-Nesbet perturbation theory32. In the original QSCI work, the dimension of the subspace (i.e., the number of determinants included in the CI expansion) was rather limited, with at most around a hundred determinants, in part due to a limited variational space when using rather small basis sets.
@@
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.33, under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 33 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is essentially identical to the original SCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)34, 35, with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing36 and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers37, spin state energetics of Fe(III)-NTA 38 and methylene39, the treatment of excited states40, and inclusion of environment effects through density matrix embedding theory41.
@@
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed42. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
@@
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem43, 44.
@@
A second issue arises from the noisy nature of currently available quantum devices. Even though the target ansatz is, in principle, particle-conserving (i.e., has a desired number of $\alpha$ and $\beta$ electrons), the presence of hardware noise leads to many samples having the wrong number of particles. For example, in Ref. 33, only 0.44% of samples for the iron-sulfur cluster [2Fe-2S] (30e,20o) are within the target particle sector. Solutions to this problem have been proposed in the form of methods like the self-consistent configuration recovery (S-CORE)33, which seek to recover nearby bit-strings with a correct particle number from the set of erroneous bitstrings.
@@
In our work, we will assume that these problems could be solved by a combination of improvements in quantum hardware and algorithms. Thus, we will not consider the state preparation or the noise problem but rather assume that an exact ansatz can be found. To this end, we rely on determinants sampled from CASCI wave functions. This approach eliminates complications related to noisy quantum devices, allowing us to focus on assessing the intrinsic efficiency and the quality of QSCI expansions. Ultimately, we demonstrate that even with an ideal sampling wave function in a noise-free environment, the stochastic nature of sampling itself either introduces significant inefficiencies or leads to less compact wave functions. Less compact CI expansions (i.e., more determinants for the same energy) lead to higher (classical) computational costs associated with the diagonalization of the CI Hamiltonian for the same final accuracy in the computation. These limitations undermine any practical advantage of QSCI over well-established classical SCI heuristics. Furthermore, enhancements like ext-SQD40, which add determinants post-sampling, could equally be applied to classical SCI methods, negating any unique benefit of quantum sampling.
2 Theory
@@
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy45, which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions10, 11, 12, 13, a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contributes significantly to the electronic wave function for most molecular systems 46.
@@
In addition to QSCI, we will also consider a representative of classical SCI. In particular, we will consider Heat-bath Configuration Interaction12 (HCI). In HCI, important determinants for the CI expansion are added based on their expected contribution to the wave function. It uses a parameter $\varepsilon$ to screen a potential determinant $\left|\Phi_{I}\right\rangle$ and includes it if any of the existing determinants in the SCI expansion satisfy
@@
The HCI method works in an iterative fashion. Starting from some initial CI expansion (e.g., the HF determinant), the following steps are repeated: 1) the CI Hamiltonian is diagonalized, and the CI coefficients of the current determinant set are obtained; 2) determinants are added according to the criterion in eq. (3), with $\varepsilon$ as a parameter. These steps are repeated until no more determinants can be added according to eq. (3) (in the original work12, when the number of new determinants is less than 1% of the number of determinants already selected).
3 Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ47 basis) of the nitrogen molecule, were obtained using PySCF 48. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan49, 50. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library51 was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)34 and local UCJ (LUCJ)35 ansatz using ffsim52. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware33.
4 Results and Discussion
@@
4.1 Potential energy curves of the nitrogen molecule
@@
Figure 1: Ground-state potential energy curve of N2/cc-pVDZ . The left panel shows total energies from HCI (solid lines), QSCI (dots), and CASCI (solid black line) wave functions with different thresholds $\varepsilon$ (HCI) and number of samples $N$ (QSCI). The middle panel shows energy errors relative to the CASCI reference. The right panel shows the number of determinants in the CI expansions. For the QSCI results, we include ten independent runs and indicate the variance across runs with error bars at a 95% confidence interval.
@@
where $\alpha$ is a scaling parameter and $p_{I}=|c_{I}|^{2}$ is the original CI vector probability of a given determinant $\left|\Phi_{I}\right\rangle$. Such a probability distribution allows us to tune the sampling by changing the weights of the determinants in the original CI vector, making it more likely (or less likely) to find new determinants depending on the value of $\alpha$. The original distribution of the determinants is recovered with $\alpha=1$. Letting $\alpha>1$ emphasizes the larger weights in the probability distribution, while letting $\alpha<1$ emphasizes smaller weights in the probability distribution. At $\alpha=0$, this distribution corresponds to uniform sampling from the correct particle sector (see also Figure S1).
@@
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ34 and heavy-hex 2-LUCJ35 wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations33, although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S3. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. However, to the best of our knowledge, implementing this in practice proves to be a significant challenge and would still suffer the fundamental problems of QSCI reported above.
4.2 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan49, 50. We note that the same system has been studied in several QSCI works33, 40. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.

### v2

@@ Abstract
Quantum Selected Configuration Interaction (QSCI) methods (also known as Sample-based Quantum Diagonalization, SQD) have emerged as promising near-term approaches to solving the electronic Schrödinger equation with quantum computers. In this work, we show that QSCI methods face fundamental limitations that severely hinder their practical applicability in chemistry. Using the nitrogen molecule and the iron-sulfur cluster [2Fe-2S] as examples, we demonstrate that while QSCI can, in principle, yield high-quality CI expansions similar to classical SCI heuristics in some cases, the method struggles with inefficiencies in finding new determinants as sampling repeatedly selects already seen configurations. This inefficiency becomes especially pronounced when targeting high-accuracy results or sampling from an approximate ansatz. In cases where the sampling problem is not present, the resulting CI expansions are less compact than those generated from classical heuristics, rendering QSCI an overall more expensive method. Our findings suggest a fatal flaw in QSCI methods as the inescapable trade-off between finding sufficiently many determinants and generating compact, accurate CI expansions. This ultimately hinders utility in quantum chemistry applications as QSCI falls behind more efficient classical counterparts.
I Introduction
@@
The full configuration interaction (FCI) method provides the exact solution to the electronic Schrödinger equation (within a given basis) and is thus considered the ground truth within quantum chemistry[1, 2, 3]. However, due to the exponential scaling in computing and memory requirements, exact FCI is limited to very small molecules. One of the largest demonstrations to date was by Gao et al. on propane (C3H8)/STO-3G[4], who used a distributed FCI implementation on 256 nodes to treat a CI expansion of 1.3 trillion determinants. As a technical achievement, such work is certainly impressive. If one is willing to settle for near-exact FCI results (within, e.g., micro-Hartree precision), approximate methods such as selected configuration interaction (SCI) can obtain results much more affordably[5]. Various developments within electronic structure theory have focused on achieving near-FCI accuracy at a lower cost[3], including methods based on the density matrix renormalization group (DMRG)[6], many-body expanded FCI (MBE-FCI)[7, 8, 9], and various flavors of SCI[10, 11, 12, 13]. Recently, such methods were put to the test in a “blind challenge” of predicting the ground-state energy of the benzene molecule (C6H6) in a cc-pVDZ basis set[2]. It revealed that a firm consensus between various near-exact FCI methods has still not been reached, even when agreement from calculations on smaller systems would have suggested so.
@@
Applying quantum computers within the field of quantum chemistry has received significant attention in recent years[14, 15, 16, 17]. This is partly due to the expectation that quantum computers should be a natural candidate for solving the electronic Schrödinger equation as they inherently operate within an exponentially large Hilbert space. In the long term, provided fault-tolerant devices become a reality, quantum computers could, for example, be used to obtain ground state energies using quantum phase estimation[18, 19]. However, current near-term quantum computers are noisy and rather limited in terms of qubit count, error rates, and execution speed, which limits their practical advantage over classical computing architectures[20]. This has motivated the development of hybrid algorithms that take advantage of both quantum and classical computing architectures, resulting in shallower quantum circuits, which are more realistic options for practical quantum computing applications in the near term[21, 22, 23]. The variational quantum eigensolver (VQE) algorithm[24, 25, 26] has emerged as a particularly popular near-term method. VQE-type methods rely on a parameterized ansatz, which is iteratively optimized using a hybrid quantum-classical feedback loop. The quantum device prepares a state and measures the expectation values over Pauli operators. The classical computer collects these measurements to evaluate expectation values over the fermionic Hamiltonian and updates the parameters of the parameterized ansatz using classical optimization methods. Although popular, VQE-type methods have several serious drawbacks related to the measurement process[27] (which inherently has statistical noise) and the expressibility of the ansatz[28, 29, 26].
@@
The quantum selected configuration interaction (QSCI) method, introduced by Kanno et al.[30, 31], addresses some of the challenges faced with VQE by drastically reducing the quantum workload, leveraging quantum devices solely for determinant selection. The central idea of the method is that some approximate ground state is prepared on a quantum computer, which is then used to draw samples corresponding to determinants in the computational basis. The CI Hamiltonian is then classically diagonalized in the basis of the selected determinants using conventional SCI diagonalization routines. We note that the classic diagonalization eventually becomes the computational bottleneck of the algorithm, provided enough determinants are discovered. Thus, the “quantum” part of QSCI lies solely in the selection of determinants for the SCI expansion. Such a division of work between quantum and classical resources has some clear advantages. First, since the quantum device is only used for selection, while energies are obtained classically, one removes the statistical errors associated with Hamiltonian measurement. Second, the CI vector is accessible on the classical device after the diagonalization, which allows computing expectation values over other quantum mechanical operators or correcting the energy of the variational wave function with Epstein-Nesbet perturbation theory[32]. In the original QSCI work, the dimension of the subspace (i.e., the number of determinants included in the CI expansion) was rather limited, with at most around a hundred determinants, in part due to a limited variational space when using rather small basis sets.
@@
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.[33], under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 33 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is essentially identical to the original SCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)[34, 35], with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing[36] and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers[37], spin state energetics of Fe(III)-NTA [38] and methylene[39], the treatment of excited states[40], inclusion of environment effects through density matrix embedding theory[41], and in combination with implicit-solvation models[42].
@@
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed[43]. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
@@
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem[44, 45]. Later, similar developments under the SQD moniker have also appeared[46].
@@
A second issue arises from the noisy nature of currently available quantum devices. Even though the target ansatz is, in principle, particle-conserving (i.e., has a desired number of $\alpha$ and $\beta$ electrons), the presence of hardware noise leads to many samples having the wrong number of particles. For example, in Ref. 33, only 0.44% of samples for the iron-sulfur cluster [2Fe-2S] (30e,20o) are within the target particle sector. Solutions to this problem have been proposed in the form of methods like the self-consistent configuration recovery (S-CORE)[33], which seek to recover nearby bit-strings with a correct particle number from the set of erroneous bitstrings.
@@
In our work, we will assume that these problems could be solved by a combination of improvements in quantum hardware and algorithms. Thus, we will not consider the state preparation or the noise problem but rather assume that an exact ansatz can be found. To this end, we rely on determinants sampled from CASCI wave functions. This approach eliminates complications related to noisy quantum devices, allowing us to focus on assessing the intrinsic efficiency and the quality of QSCI expansions. Ultimately, we demonstrate that even with an ideal sampling wave function in a noise-free environment, the stochastic nature of sampling itself either introduces significant inefficiencies or leads to less compact wave functions. Less compact CI expansions (i.e., more determinants for the same energy) lead to higher (classical) computational costs associated with the diagonalization of the CI Hamiltonian for the same final accuracy in the computation. These limitations undermine any practical advantage of QSCI over well-established classical SCI heuristics. Furthermore, enhancements like ext-SQD[40], which add determinants post-sampling, could equally be applied to classical SCI methods, negating any unique benefit of quantum sampling.
II Theory
@@
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy[47], which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions[10, 11, 12, 13], a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contributes significantly to the electronic wave function for most molecular systems [48].
@@
In addition to QSCI, we will also consider a representative of classical SCI. In particular, we will consider Heat-bath Configuration Interaction[12] (HCI). In HCI, important determinants for the CI expansion are added based on their expected contribution to the wave function. It uses a parameter $\varepsilon$ to screen a potential determinant $\left|\Phi_{I}\right\rangle$ and includes it if any of the existing determinants in the SCI expansion satisfy
@@
The HCI method works in an iterative fashion. Starting from some initial CI expansion (e.g., the HF determinant), the following steps are repeated: 1) the CI Hamiltonian is diagonalized, and the CI coefficients of the current determinant set are obtained; 2) determinants are added according to the criterion in eq. (3), with $\varepsilon$ as a parameter. These steps are repeated until no more determinants can be added according to eq. (3) (in the original work[12], when the number of new determinants is less than 1% of the number of determinants already selected).
III Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ[49] basis) of the nitrogen molecule, were obtained using PySCF [50]. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan[51, 52]. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library[53] was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)[34] and local UCJ (LUCJ)[35] ansatz using ffsim[54]. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware[33].
IV Results and Discussion
@@
IV.1 Potential energy curves of the nitrogen molecule
@@
Figure 1: Ground-state potential energy curve of N2/cc-pVDZ (10e,22o). The left panel shows total energies from HCI (solid lines), QSCI (dots), and CASCI (solid black line) wave functions with different thresholds $\varepsilon$ (HCI) and number of samples $N$ (QSCI). The middle panel shows energy errors relative to the CASCI reference. The right panel shows the number of determinants in the CI expansions. For the QSCI results, we include ten independent runs and indicate the variance across runs with error bars at a 95% confidence interval.
@@ IV.2 Scaled probability distributions
where $\alpha$ is a scaling parameter and $p_{I}=|c_{I}|^{2}$ is the original CI vector probability of a given determinant $\left|\Phi_{I}\right\rangle$. This can be seen as maintaining the support (configuration space) of the FCI wave function but changing its weight distribution. Such a probability distribution allows us to tune the sampling by changing the weights of the determinants in the original CI vector, making it more likely (or less likely) to find new determinants depending on the value of $\alpha$. The original distribution of the determinants is recovered with $\alpha=1$. Letting $\alpha>1$ emphasizes the larger weights in the probability distribution, while letting $\alpha<1$ emphasizes smaller weights in the probability distribution. At $\alpha=0$, this distribution corresponds to uniform sampling from the correct particle sector (see also Figure S1).
@@
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ[34] and heavy-hex 2-LUCJ[35] wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations[33], although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S3. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. However, to the best of our knowledge, implementing this in practice proves to be a significant challenge and would still suffer the fundamental problems of QSCI reported above.
IV.3 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan[51, 52]. We note that the same system has been studied in several QSCI works[33, 40]. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.